Determine wether the lines through the pairs of points are parallel,perpendiular, or neither.
12x+4y = 16
27x+9y = 40

Answer by checkley71(8403)

12X+4Y=16 OR 4Y=-12X+16 OR Y=-12X/4+16/4 OR Y=-3X+4
27X+9Y=40 OR 9Y=-27X+40 OR Y=-27X/9+40/9 OR Y=-3X+40/9
THESE 2 EQUATIONS HAVE THE SAME SLOPE (-3) THEREFORE THEY ARE PARALLEL LINES.